Michel’s class had 4 A’s, 10 B’s, 13 C’s, 8 D’s, and 2 F’s.
(a) What is the ratio of A’s to B’s?
(b) What is the ratio of F’s to total students?

Answer by drk(1908)

(a) What is the ratio of A’s to B’s?
There are 4 As and 10 B's, so the ratio of A / B = 4/10 = 2/5.
--
(b) What is the ratio of F’s to total students?
There are 2 F's and 37 total students, so ratio of F / total = 2/37.